OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3. y = 3x - 14 y = one thirdx + 2 y = -3x + 22 y = -one thirdx + 8

OpenStudy (anonymous):

Parallel to : \(y = \frac{1}{3}x - 3\) So the slope of this line is \(\large \frac{1}{3}\) So, m = 1/3 Use the equation: \[\large y-y_1 = m(x-x_1)\] \[y - 4 = \frac{1}{3} ( x - 6)\] \[\large y - 4 = \frac{1}{3}x - 2\] Add 4 both the sides and find the equation..